a driver traveling at 12m/s sees a stalled vehicle 17m ahead of his car. he steps on the brakes to slowdown the car. At what rate must the car slow down so that it stops just before hitting the stalled vehicle? Assume that the drivers reaction time is 0.25s. Reaction time is the time for a person to notice , think, and act in response to a given situation.
1
Expert's answer
2012-08-10T11:02:16-0400
The driver steps the brakes after 0.25s after he sees a stalled vehicle. During that time his car goes 0.25s*12m/s = 3m so the distance between the car and the vehicle becoms 17m-3m = 14m. So, the maximum stopping distance is 14m. To stop the car that moves at 12m/s within 14m the deceleration time must be

t = 14m/((12m/s)/2) = 2.(3)s,

and so deceleration needed is

a = (12m/s) / 2.(3)s ≈ 5.1429 m/s².